# Privacy-Preserving Federated Learning with Differentially Private Hyperdimensional Computing

### v2

@@ Abstract
Federated Learning (FL) is essential for efficient data exchange in Internet of Things (IoT) environments, as it trains Machine Learning (ML) models locally and shares only model updates. However, FL is vulnerable to privacy threats like model inversion and membership inference attacks, which can expose sensitive training data. To address these privacy concerns, Differential Privacy (DP) mechanisms are often applied. Yet, adding DP noise to black-box ML models degrades performance, especially in dynamic IoT systems where continuous, lifelong FL learning accumulates excessive noise over time. To mitigate this issue, we introduce Federated HyperDimensional computing with Privacy-preserving (FedHDPrivacy), an eXplainable Artificial Intelligence (XAI) framework that combines the neuro-symbolic paradigm with DP. FedHDPrivacy carefully manages the balance between privacy and performance by theoretically tracking cumulative noise from previous rounds and adding only the necessary incremental noise to meet privacy requirements. In a real-world case study involving in-process monitoring of manufacturing machining operations, FedHDPrivacy demonstrates robust performance, outperforming standard FL frameworks-including Federated Averaging (FedAvg), Federated Stochastic Gradient Descent (FedSGD), Federated Proximal (FedProx), Federated Normalized Averaging (FedNova), and Federated Adam (FedAdam)-by up to $38\%$. FedHDPrivacy also shows potential for future enhancements, such as multimodal data fusion.
@@ 1 Introduction
Differential Privacy (DP) has emerged as a more practical and robust solution for safeguarding privacy in AI systems. When applied to an ML model, DP protects against model inversion and membership inference attacks, making it a suitable approach for securing FL frameworks. Recent advancements have enhanced the viability of DP, with companies like Google, Microsoft, and Apple successfully integrating DP into their data collection practices to improve user privacy while retaining data utility [21, 22, 23]. DP achieves this by adding carefully calibrated noise to the data or model, ensuring that individual records remain private during analysis [24, 25]. This noise addition ensures that the inclusion or exclusion of any single data point does not significantly impact the overall analysis, thereby preserving individual privacy. As illustrated in Figure 2, DP can effectively neutralize model inversion attacks, where adversaries attempting to reconstruct training samples by querying the model will only retrieve random, non-informative signals unrelated to the actual training data. Additionally, DP can thwart membership inference attacks by preventing adversaries from determining whether a particular sample was part of the training set.
DP’s ability to protect sensitive information while preserving data utility makes it a preferred choice among privacy-preserving techniques. In a FL structure, as depicted in Figure 2, DP can be applied to local models before they are transmitted to the central server. As a result, these secured local models contribute to the global model, which is formed by aggregating the secured local models, thereby maintaining the security of the global model as well. The level of privacy can also be adjusted at the server level, ensuring that the local models remain secure and resistant to potential breaches by an untrusted server. This configuration protects the global model from server-side attacks, making it difficult for adversaries to compromise the server. Additionally, the communication between the server and clients is secure because the updates exchanged are already secured, rendering eavesdropping attacks ineffective. Furthermore, since the server is secure, a malicious participant cannot extract any information from the secured global model through interaction with the server. In conclusion, while anonymization, HE, and SMC each offer specific advantages for privacy protection in FL, DP stands out as the most effective and practical solution, providing a strong balance between privacy and usability in real-world applications.
@@
While DP is highly effective at safeguarding sensitive information and maintaining data utility, it also presents challenges related to balancing privacy and accuracy. In the context of FL combined with DP, the amount of noise introduced must be carefully managed to balance privacy and accuracy. Excessive noise can degrade model performance, while insufficient noise can lead to privacy breaches. Therefore, evaluating privacy leakage for a given privacy budget is crucial before deploying models or releasing datasets [12]. Moreover, ML models often function as complex black-box systems, where their internal mechanisms are not easily understood or transparent. This lack of transparency is particularly problematic in safety-critical fields, where the traceability and explainability of decision-making processes are crucial [26]. Additionally, the secure handling of data is essential, especially when dealing with sensitive and personal information. In the FL framework, where data privacy is paramount, DP is widely adopted to protect sensitive information by adding noise to the data or model parameters. However, implementing DP within FL presents a significant challenge: as noise is added in each training round to maintain privacy, the cumulative effect can lead to substantial degradation in model performance. This issue is exacerbated by the black-box nature of ML models, making it difficult to determine the optimal amount of noise to add without compromising the model’s accuracy [27].
@@
We introduce Federated HyperDimensional computing with Privacy-preserving (FedHDPrivacy), an eXplainable Artificial Intelligence (XAI) framework for FL, to address the privacy challenges. This framework combines DP and HyperDimensional computing (HD) to ensure both security and accuracy during model aggregation. By integrating XAI techniques, FedHDPrivacy allows for precise calculation of the noise required in each round, tracking the cumulative noise added in previous rounds and adjusting the noise accordingly in subsequent rounds. This approach ensures that only the minimal necessary noise is introduced, effectively balancing privacy and accuracy. The FedHDPrivacy framework offers a robust solution by controlling the noise added to both the clients’ and the server’s models in each round, thereby avoiding the degradation of model performance due to excessive noise. This secure and explainable FL framework not only safeguards confidential information but also preserves the utility and efficiency of the AI models involved. Moreover, our framework is specifically designed to address the challenges of continuous learning in IoT systems under the FL paradigm. It ensures that the model remains up-to-date and effective in dynamic environments, thereby enhancing the practical utility of FL in real-world IoT scenarios. Our contributions in this paper are summarized as follows:
@@
    We introduce FedHDPrivacy, an explainable framework that enhances transparency and understanding in the interactions between ML models and DP within an FL structure, ensuring both security and interpretability.
@@
    Our framework accurately determines the necessary noise levels to secure both client and central server models, while preventing the overestimation of DP noise by computing the cumulative noise from previous rounds.
@@
    FedHDPrivacy ensures an optimal balance between privacy protection and model accuracy by avoiding excessive noise addition throughout the iterative training process.
The remainder of this paper is structured as follows: Section 2 reviews existing FL frameworks, examines DP as a mechanism for safeguarding ML models in FL, and explores the integration of HD with FL and DP. In Section 3, we define key concepts related to HD and DP. Section 4 outlines the proposed FedHDPrivacy framework and its implementation. Section 5 details the experimental setup used in a real-world IoT scenario. The results of the experiments are presented and analyzed in Section 6. Finally, Section 7 summarizes the contributions of this work in developing privacy-preserving FL models for IoT and offers suggestions for future research directions.
@@ 2.1 Federated Learning
FL offers significant advantages for IoT applications by enabling local data processing on devices while only transmitting model updates to a central server. This approach minimizes the risk of exposing sensitive information to third parties, ensuring compliance with strict data privacy regulations, such as the general data protection regulation [28]. Consequently, FL is well-suited for secure IoT systems. Moreover, by leveraging the computational power and diverse datasets available across various IoT devices, FL can speed up the training process and achieve higher accuracy compared to centralized AI methods, which often face challenges due to limited data and resources [29]. Additionally, the decentralized structure of FL improves the scalability of intelligent networks.
@@
FL has progressed through the development of various frameworks aimed at addressing the challenges of distributed optimization and the diversity of client data and resources. Federated Stochastic Gradient Descent (FedSGD) and Federated Averaging (FedAvg), both introduced by McMahan et al. [8], offer foundational approaches for FL. FedSGD focuses on aggregating gradients from multiple clients, while FedAvg aggregates model parameters by averaging them across clients, providing a straightforward yet effective method for distributed learning. Federated Proximal (FedProx), developed by Li et al. [30], enhances FedAvg by incorporating a proximal term into the local objective functions, which mitigates the effects of client heterogeneity and improves convergence. Federated Normalized Averaging (FedNova), introduced by Wang et al. [31], addresses the issue of objective inconsistency by normalizing client updates, ensuring a more robust and consistent aggregation process. Additionally, Federated Adam (FedAdam), proposed by Reddi et al. [32], extends the Adam optimizer to the federated setting. FedAdam adapts the learning rates for each client by leveraging second-order moment estimates, thus offering better convergence properties, especially in non-identically distributed data scenarios. These FL frameworks facilitate the distribution of the learning process among clients without transmitting raw data, thereby safeguarding confidential information. However, they remain vulnerable to model inversion and membership inference attacks, underscoring the need for DP to secure these frameworks against such threats.
2.2 Differential Privacy
Several studies have explored the integration of DP within FL frameworks to enhance system security. Gong et al. proposed a framework that combines DP with HE to protect client gradients from the central server, although challenges remain in fully preventing information leakage even with the application of DP [33]. Wu et al. applied DP to multi-task learning in FL, perturbing parameters with Gaussian noise to maintain privacy, yet their method remains vulnerable to certain attacks, such as label-flipping and model inversion [34]. Cyffers et al. introduced a decentralized FL approach using a relaxed version of local DP, enabling data analysis across devices while balancing privacy and utility [24]. However, the decentralized nature of this protocol leaves it open to data poisoning and other attacks [35]. Tran et al. developed a secure decentralized training framework that eliminates the need for a central server by rotating the master node role among clients [36], but this approach still struggles to fully protect against collusion attacks [35]. Zhao et al. proposed the sharing of partial gradient parameters combined with Gaussian noise and introduced an additional proxy to enhance client anonymity [37]. While this method strengthens privacy, it requires further evaluation against inference attacks. Yin et al. suggested a privacy-preserving strategy using functional encryption and Bayesian DP, which secures client-server communication and monitors privacy leakage [38]. However, this method relies on a trusted third party, introducing the risk of model inversion attacks.
Due to the dynamic nature of IoT, FL must support continuous training over time, adapting to newly generated data in each communication round. While DP secures the training samples by adding noise in each round, the cumulative effect of this noise can eventually degrade model accuracy. To address this issue, an XAI model is required to monitor and manage the cumulative noise from previous rounds. By accurately tracking this noise, the XAI model can ensure that only the necessary additional noise is introduced, calculated as the difference between the required noise and the cumulative noise from earlier rounds. This approach helps maintain a balance between privacy and accuracy over the lifetime of the FL process.
2.3 Hyperdimensional Computing
HD takes significant inspiration from the architecture of the human brain, utilizing high-dimensional vectors to simulate cognitive functions [39, 40, 41]. This advanced approach allows HD to mimic brain-like methods for processing, analyzing, and storing information across a variety of cognitive tasks [42, 43, 44]. Several studies have explored the application of HD within FL frameworks. Zhao et al. [45] developed an FL framework using HD, demonstrating the energy and communication efficiency of their model. While they highlighted the robustness of HD against noise, they did not address the privacy implications. Li et al. [46] introduced an HD-based FL framework aimed at reducing the computational and communication overhead associated with local training and model updates. Hsieh et al. [47] bipolarized model parameters to decrease communication costs in an HD-based FL framework, yet they did not provide specific defenses against model inversion and membership inference attacks. As an XAI model, HD is susceptible to model inversion and inference attacks [48], making it imperative to integrate privacy-preserving mechanisms to protect FL systems based on HD from adversarial threats.
Some studies have explored privacy-preserving techniques for HD models. Hernandez-Cano et al. [49] implemented intelligent noise injection that identifies and randomizes insignificant features of the model in the original space. However, their work does not clearly define how much noise is necessary to achieve a specific level of privacy. Khaleghi et al. [50] introduced DP to HD models to enhance security, but they did not predict the impact of noise on the model’s accuracy. Jalil Piran et al. [48] investigated the transparency of combining DP with HD models and predicted the effect of noise on HD’s performance at various privacy levels. Nonetheless, they did not consider these aspects within a FL structure. Therefore, it is essential to explore the transparency and explainability of HD models in an FL context, allowing for the precise addition of noise to both client models and the global model. This approach will help secure training data while minimizing the impact on model accuracy.
3 Preliminary
3.1 Hyperdimensional Computing
3.1.1 Encoding
The encoding phase is essential for transforming input data into hypervectors, which form the foundation of the high-dimensional computational framework. This process ensures that every component of a hypervector plays an equal role in representing the encoded information, thereby maintaining a balanced and comprehensive data representation. An input sample, denoted as x, is mapped into a hypervector $\vec{H}$ as shown in Equation (1). This mapping follows the principles of the random vector functional link model, a technique that generates random projections to map input features into a higher-dimensional space [55, 56]. This model enhances the representational capacity by introducing non-linearities, ensuring that the transformed hypervectors effectively capture the complexity of the input data. After this transformation, the elements of the hypervector are binarized.
$\phi:\mathbf{x}\rightarrow\vec{H}$ (1)
3.1.2 Training
In the training phase, class hypervectors, denoted as $\vec{C}^{s}$, are created for each class $s$, where $s$ ranges from $1$ to $S$ (the total number of classes). These hypervectors are formed by summing all associated hypervectors for each class, as described in Equation (2). Here, $\vec{H}^{s}$ represents the hypervectors linked to class $s$. This summation combines the features of all training samples within a class into a single high-dimensional representation, crucial for accurate classification. The aggregation process captures the shared characteristics of each class, enabling the HD model to effectively distinguish between different classes during inference. Creating class hypervectors is a key step in the HD training process, laying the groundwork for accurate pattern recognition and prediction capabilities.
$\vec{C}^{s}=\sum\vec{H}^{s}$ (2)
3.1.3 Inference
3.1.4 Retraining
Retraining plays a crucial role in enhancing the accuracy and adaptability of the HD model, particularly in continuous learning environments like FL. Instead of retraining the model from scratch with all previous data, retraining in HD focuses on correcting misclassifications by fine-tuning the model using new training samples. This process involves comparing the hypervectors of new data points with the existing class hypervectors. If a hypervector, $\vec{H}^{s}$, is mistakenly classified into an incorrect class $s^{{}^{\prime}}$ rather than its true class $s$, the model is updated to correct this error. The class hypervectors are adjusted by adding the hypervector to its correct class and subtracting it from the incorrect one, as expressed in the following equations:
$\vec{C}^{s}$ $\leftarrow\vec{C}^{s}+\vec{H}^{s}$ (4)
$\vec{C}^{s^{{}^{\prime}}}$ $\leftarrow\vec{C}^{s^{{}^{\prime}}}-\vec{H}^{s}$
3.2 Differential Privacy
In this work, we address the vital challenge of protecting privacy within FL frameworks for IoT systems. A key aspect of this protection involves defending against model inversion and membership inference attacks, which can expose sensitive information. Traditional systems often inadvertently compromise data privacy due to the broad access they require to raw data. To mitigate these risks, we employ DP techniques. DP mechanisms ensure that only privacy-preserving outputs, modified to adhere to DP principles, are exposed, rather than the raw data itself. This strategy significantly enhances security, reducing the likelihood of privacy breaches and fortifying the AI models against potential security threats. The variables used throughout this work are summarized in Table 1.
@@ 3.2 Differential Privacy
DP strengthens the security of AI models by restricting access to raw data and allowing only modified outputs generated through DP techniques. DP achieves this by intentionally adding noise to the data, which helps to mask the original information. This process is controlled by two important parameters: the privacy budget ($\epsilon$) and the privacy loss threshold ($\delta$). These parameters determine the amount of noise added and the level of privacy protection, significantly reducing the chances of identifying specific data points. DP ensures that the privacy of individuals is safeguarded while still allowing the data to be useful for analysis, providing a powerful approach to protecting sensitive information in FL environments against threats such as model inversion and membership inference attacks.

### v3

@@ Abstract
Federated Learning (FL) has become a key method for preserving data privacy in Internet of Things (IoT) environments, as it trains Machine Learning (ML) models locally while transmitting only model updates. Despite this design, FL remains susceptible to threats such as model inversion and membership inference attacks, which can reveal private training data. Differential Privacy (DP) techniques are often introduced to mitigate these risks, but simply injecting DP noise into black-box ML models can compromise accuracy, particularly in dynamic IoT contexts, where continuous, lifelong learning leads to excessive noise accumulation. To address this challenge, we propose Federated HyperDimensional computing with Privacy-preserving (FedHDPrivacy), an eXplainable Artificial Intelligence (XAI) framework that integrates neuro-symbolic computing and DP. Unlike conventional approaches, FedHDPrivacy actively monitors the cumulative noise across learning rounds and adds only the additional noise required to satisfy privacy constraints. In a real-world application for monitoring manufacturing machining processes, FedHDPrivacy maintains high performance while surpassing standard FL frameworks — Federated Averaging (FedAvg), Federated Proximal (FedProx), Federated Normalized Averaging (FedNova), and Federated Optimization (FedOpt) — by up to 37%. Looking ahead, FedHDPrivacy offers a promising avenue for further enhancements, such as incorporating multimodal data fusion.
@@ 1 Introduction
Differential Privacy (DP) has emerged as a practical and effective solution for safeguarding privacy in AI systems, particularly within FL frameworks. By introducing carefully calibrated noise to data or models, DP ensures that individual records remain private during analysis, mitigating risks from model inversion, and membership inference attacks [21, 22]. This approach has been successfully implemented by companies like Google, Microsoft, and Apple to enhance user privacy while preserving data utility [23, 24, 25]. In FL structures, as depicted in Figure 2, DP can be applied to local models before transmission to the central server, ensuring that both local and global models remain secure against potential breaches, even in scenarios involving untrusted servers. This configuration also safeguards communication channels, rendering eavesdropping and malicious participant attacks ineffective. Compared to methods like anonymization, HE, and SMC, DP strikes an optimal balance between privacy and practicality, making it a preferred choice for real-world applications.
@@
While DP is highly effective at safeguarding sensitive information and maintaining data utility, it presents challenges in balancing privacy and accuracy [26]. In the context of continuous FL combined with DP, noise is added during each communication round to protect privacy, but this cumulative noise can significantly degrade model performance over time. Excessive noise reduces accuracy, while insufficient noise increases the risk of privacy breaches. To address this, it is crucial to develop privacy-preserving mechanisms capable of tracking the noise accumulation across communication rounds and adding only the necessary noise in each round to guarantee privacy without unnecessarily compromising accuracy. Moreover, evaluating privacy leakage for a given privacy budget is essential before deploying models or releasing datasets [12]. This challenge is further complicated by the black-box nature of Machine Learning (ML) models, where their internal processes lack transparency [27]. In safety-critical applications, where both explainability and privacy are critical, this lack of traceability becomes a significant issue. Therefore, the development of adaptive noise management strategies within FL frameworks is essential to maintain a balance between privacy and accuracy, ensuring the effectiveness of privacy-preserving mechanisms without undermining model performance [28].
@@
We propose Federated HyperDimensional computing with Privacy-preserving (FedHDPrivacy), an eXplainable Artificial Intelligence (XAI) framework for FL that addresses the privacy challenges in dynamic IoT environments. By combining DP and HyperDimensional computing (HD), FedHDPrivacy ensures robust privacy protection while maintaining high model accuracy during the aggregation process. The framework leverages XAI techniques to dynamically track the cumulative noise added in previous communication rounds and adjust the noise in subsequent rounds, introducing only the minimal necessary noise to guarantee privacy without compromising accuracy. This adaptive approach prevents excessive noise accumulation that could degrade model performance, securing both client and server models at every training round. Furthermore, FedHDPrivacy is designed to support continuous learning in IoT systems, enabling the global model to remain effective and responsive in dynamic environments. By safeguarding sensitive information and preserving model utility, FedHDPrivacy offers a secure, efficient, and explainable solution tailored to the real-world challenges of FL in IoT scenarios. Our contributions in this paper are summarized as follows:
@@
    We introduce FedHDPrivacy, a novel and explainable framework that integrates HD with DP within a FL structure, ensuring robust privacy protection and model interpretability.
@@
    Our framework accurately tracks cumulative noise from previous communication rounds and determines the incremental noise required, ensuring compliance with DP guarantees while avoiding unnecessary noise addition.
@@
    FedHDPrivacy demonstrates adaptability to dynamic IoT environments by enabling continuous learning without performance degradation, effectively balancing privacy, accuracy, and efficiency.
  4. 4.
    The proposed framework outperforms existing FL methods — Federated Averaging (FedAvg), Federated Proximal (FedProx), Federated Normalized Averaging (FedNova), and Federated Optimization (FedOpt) — in accuracy (up to 37%) and computational efficiency, as validated through extensive experiments using real-world IoT manufacturing data.
  5. 5.
    We provide a comprehensive analysis of hyperparameter effects, including hypervector size and privacy budget, on model performance, training time, and energy efficiency, offering insights into practical implementations of privacy-preserving FL systems.
The remainder of this paper is structured as follows: Section 2 reviews existing FL frameworks, examines DP as a mechanism for safeguarding ML models in FL, and explores the integration of HD with FL and DP. In Section 3, we define key concepts related to HD and DP. Section 4 outlines the proposed FedHDPrivacy framework and its implementation. Section 5 details the experimental setup used in a real-world IoT scenario. The results of the experiments are presented and analyzed in Section 6. Finally, Section 7 summarizes the contributions of this work in developing privacy-preserving FL models for IoT and offers suggestions for future research directions. The variables used throughout this work are summarized in Table 1.
@@ 2.1 Federated Learning
FL offers substantial benefits for IoT applications by enabling local data processing on devices while transmitting only model updates to a central server. This decentralized approach minimizes the risk of exposing sensitive information to third parties and ensures compliance with strict data privacy regulations, such as the general data protection regulation [29]. FL is particularly well-suited for secure IoT systems, as it leverages the computational power and diverse datasets of IoT devices to accelerate training and achieve higher accuracy compared to centralized AI methods, which often struggle with limited data and resources [30]. Additionally, FL’s decentralized structure enhances scalability, making it ideal for expanding intelligent networks.
@@
The evolution of FL has seen the development of several frameworks addressing distributed optimization challenges and the heterogeneity of client data and resources. Federated Stochastic Gradient Descent (FedSGD) and FedAvg, introduced by McMahan et al.  [8], laid the foundation for FL. While FedSGD aggregates gradients from multiple clients, FedAvg simplifies the process by averaging model parameters across clients, offering an effective approach for distributed learning. To address client heterogeneity and improve convergence, Li et al.  [31] proposed FedProx, which introduces a proximal term in the local objective functions. FedNova, developed by Wang et al.  [32], resolves objective inconsistencies by normalizing client updates, ensuring a more robust and consistent aggregation process. Additionally, Federated Adam (FedAdam), introduced by Reddi et al.  [33], extends the Adam optimizer to the federated setting by adapting learning rates for each client through second-order moment estimates, improving convergence in non-identically distributed data scenarios. Furthermore, the FedOpt framework, also proposed by Reddi et al.  [33], generalizes client and server optimization strategies, allowing for the integration of different optimizers such as SGD and Adam on the server side. FedOpt provides flexibility in tuning the aggregation process, leading to better performance across diverse FL scenarios.
These FL frameworks enable distributed learning among clients without transmitting raw data, thereby safeguarding confidential information. However, they remain vulnerable to model inversion and membership inference attacks, highlighting the need for DP to secure these frameworks against such threats.
@@
2.2 Differential Privacy
Several studies have investigated the integration of DP within FL frameworks to enhance system security. Gong et al. proposed a framework that combines DP with HE to protect client gradients from the central server, though challenges remain in fully preventing information leakage even with DP [34]. Wu et al. applied DP to multi-task learning in FL by perturbing parameters with Gaussian noise to preserve privacy; however, their method remains susceptible to attacks such as label-flipping and model inversion [35]. Cyffers et al. introduced a decentralized FL approach utilizing a relaxed form of local DP, enabling data analysis across devices while balancing privacy and utility [21]. Nevertheless, the decentralized nature of this protocol leaves it vulnerable to data poisoning and other security threats [36]. Tran et al. developed a secure decentralized training framework that eliminates the need for a central server by rotating the master node role among clients [37], but this approach struggles to fully defend against collusion attacks [36]. Zhao et al. proposed a method involving the sharing of partial gradient parameters combined with Gaussian noise and introduced an additional proxy to enhance client anonymity [38]. While this technique strengthens privacy, it requires further evaluation to assess its resilience against inference attacks. Yin et al. suggested a privacy-preserving strategy using functional encryption and Bayesian DP to secure client–server communication and monitor privacy leakage [39]. However, this method relies on a trusted third party, which introduces vulnerabilities to model inversion attacks. Other works have focused on adaptive encoding methods to dynamically adjust noise addition based on input data distribution, minimizing utility loss and achieving improved accuracy [40]. DP techniques have also been proposed for preserving information in cyber–physical systems by introducing random noise to generate anonymity, though these methods can impact data quality [41]. Approaches leveraging three-way attribute decisions in IoT systems have shown potential in stabilizing datasets and improving privacy by categorizing sensitive attributes before applying DP [42]. Recent advancements have further integrated dynamic privacy budget adjustment and adaptive weighting mechanisms to address challenges posed by heterogeneous IoT data, achieving improved privacy preservation and performance [43].
Given the dynamic nature of IoT systems, FL must support continuous training over time, adapting to newly generated data in each communication round. While DP provides privacy by adding noise in every round, the cumulative effect of this noise can degrade model accuracy over time. To address this challenge, an XAI model is needed to monitor and manage the cumulative noise from previous rounds. By accurately tracking this noise, the XAI model can calculate the additional noise required for privacy as the difference between the target noise level and the cumulative noise already introduced. This adaptive approach helps maintain an optimal balance between privacy and accuracy throughout the FL process, ensuring the framework remains effective in dynamic, real-world IoT environments.
2.3 Hyperdimensional Computing
HD draws significant inspiration from the human brain, leveraging high-dimensional vectors to simulate cognitive functions [44, 45, 46]. This advanced approach enables HD to mimic brain-like methods for processing, analyzing, and storing information across diverse cognitive tasks [47, 48, 49]. Several studies have explored the integration of HD within FL frameworks. Zhao et al.  [50] proposed an FL framework utilizing HD, demonstrating the model’s energy and communication efficiency. While their work highlighted HD’s robustness against noise, it did not address the privacy implications. Li et al.  [51] introduced an HD-based FL framework designed to reduce computational and communication overhead during local training and model updates. Hsieh et al.  [52] bipolarized model parameters in an HD-based FL framework to minimize communication costs, yet their work did not incorporate specific defenses against model inversion and membership inference attacks. As an XAI model, HD is inherently susceptible to model inversion and inference attacks [53], emphasizing the need for privacy-preserving mechanisms to protect HD-based FL systems from adversarial threats.
Some studies have explored privacy-preserving techniques specifically for HD models. Hernandez-Cano et al.  [54] implemented intelligent noise injection by randomizing insignificant features in the original feature space. However, their approach does not clearly define the required noise level to achieve a specific privacy guarantee. Khaleghi et al.  [55] introduced DP into HD models to enhance security but did not predict the impact of noise on model accuracy. Piran et al.  [53] examined the transparency of combining DP with HD models and evaluated noise’s effect on HD performance at various privacy levels, though their study did not address these considerations within an FL framework. Consequently, there is a critical need to explore the transparency and explainability of HD models within the context of FL. This includes developing techniques for precisely adding noise to both client models and the global model, ensuring robust data security while minimizing the impact on model accuracy.
3 Preliminary
3.1 Hyperdimensional Computing
3.1.1 Encoding
The encoding phase is essential for transforming input data into hypervectors, which form the foundation of the high-dimensional computational framework. This process ensures that every component of a hypervector plays an equal role in representing the encoded information, thereby maintaining a balanced and comprehensive data representation. An input sample, denoted as $\mathbf{x}$, is mapped into a hypervector $\vec{H}$ as shown in (1). This mapping follows the principles of the random vector functional link model, a technique that generates random projections to map input features into a higher-dimensional space [60, 61]. This model enhances the representational capacity by introducing non-linearities, ensuring that the transformed hypervectors effectively capture the complexity of the input data. After this transformation, the elements of the hypervector are binarized.
$\phi:\mathbf{x}\rightarrow\vec{H}$ (1)
3.1.2 Training
In the training phase, class hypervectors, denoted as $\vec{C}^{s}$, are created for each class $s$, where $s$ ranges from $1$ to $S$ (the total number of classes). These hypervectors are formed by summing all associated hypervectors for each class, as described in (2). Here, $\vec{H}^{s}$ represents the hypervectors linked to class $s$. This summation combines the features of all training samples within a class into a single high-dimensional representation, crucial for accurate classification. The aggregation process captures the shared characteristics of each class, enabling the HD model to effectively distinguish between different classes during inference. Creating class hypervectors is a key step in the HD training process, laying the groundwork for accurate pattern recognition and prediction capabilities.
$\vec{C}^{s}=\sum\vec{H}^{s}$ (2)
3.1.3 Inference
3.1.4 Retraining
Retraining plays a crucial role in enhancing the accuracy and adaptability of the HD model, particularly in continuous learning environments like FL. Instead of retraining the model from scratch with all previous data, retraining in HD focuses on correcting misclassifications by fine-tuning the model using new training samples. This process involves comparing the hypervectors of new data points with the existing class hypervectors. If a hypervector, $\vec{H}^{s}$, is mistakenly classified into an incorrect class $s^{\prime}$ rather than its true class $s$, the model is updated to correct this error. The class hypervectors are adjusted by adding the hypervector to its correct class and subtracting it from the incorrect one, as expressed in the following equations:
$\vec{C}^{s}$ $\leftarrow\vec{C}^{s}+\vec{H}^{s}$ (4)
$\vec{C}^{s^{{}^{\prime}}}$ $\leftarrow\vec{C}^{s^{{}^{\prime}}}-\vec{H}^{s}$
3.2 Differential Privacy
In this work, we address the practical implementation of DP within FL frameworks, specifically tailored for IoT environments. To ensure robust privacy protection, we utilize DP mechanisms to introduce noise in a controlled manner, governed by key parameters such as the privacy budget ($\epsilon$) and the privacy loss threshold ($\delta$). The following definitions and formulations provide the foundational understanding of DP concepts and their application in this work.